OpenStudy (anonymous):

Solve. 3x2 - 11x = 4.

OpenStudy (anonymous):

this one needs the quadratic formula i think

OpenStudy (anonymous):

start with \[3x^2-11x-4=0\] then use \[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \[a=3,b=-11,c=-4\]

OpenStudy (anonymous):

lol or if you know how to factor, (which i don't) it turns out that \[3x^2-11x-4=(x-4)(3x+1)\]

OpenStudy (anonymous):

then solve two equations \[x-4=0\] and also solve \[3x+1=0\]

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

you really keeping me in suspense ...

OpenStudy (anonymous):

sorry i lost all internet, and ohh its just quadratic equation?? @satellite73

OpenStudy (anonymous):

so x=4 and -1/3 @satellite73

OpenStudy (anonymous):

oui

OpenStudy (anonymous):

"just" a quadratic equation? lol yeah

OpenStudy (anonymous):

i guess i shouldn't say that.. it isn't exactly "just" a quadratic equation aha.. i cant understand it, thanks again!! @satellite73

OpenStudy (anonymous):

yw